OpenStudy (anonymous):

Write (-x-3)(-2x-3) in standard form

OpenStudy (ghazi):

use law of distribution

OpenStudy (ghazi):

\[2x^2+9x+9\]

OpenStudy (anonymous):

thats what i got

OpenStudy (ghazi):

cool :)

OpenStudy (anonymous):

Sooo how would i find the x-intercepts

OpenStudy (ghazi):

it's a quadratic equation solve for x

OpenStudy (anonymous):

How though? My math teacher is quite the retrice

OpenStudy (ghazi):

\[2x^2+6x+3x+9=0\]

OpenStudy (ghazi):

can you do it now?

OpenStudy (anonymous):

Could you do it for me so I can double check my answer?

OpenStudy (ghazi):

what you got?

OpenStudy (ghazi):

\[2x(x+3)+3(x+3)= (2x+3)(x+3)=0\]

OpenStudy (anonymous):

Yeah I'm lost...

OpenStudy (ghazi):

just use method of factorization and then you'll get x=-3/2 and -3

OpenStudy (anonymous):

okay thanks

OpenStudy (ghazi):

:) YW